# Joint Beamforming and Position Optimization for Fluid RIS-aided ISAC Systems

### v2

@@ II-A3 Channel Model
As stated before, the steering vector of the fRIS is dependent on the positions of the elements $\mathbf{p}$ as well as the azimuth angle $\phi$ and the elevate angle $\psi$. Specifically, the steering vector of the fRIS can be expressed as [25]:
$\!\!\mathbf{a}(\phi,\psi,\mathbf{p})\!=\!\begin{bmatrix}e^{\jmath\frac{2\pi}{\lambda}d_{1}(\phi,\psi,\mathbf{p}_{1})}&\!\!\!\cdots\!\!\!&e^{\jmath\frac{2\pi}{\lambda}d_{N}(\phi,\psi,\mathbf{p}_{N})}\end{bmatrix}^{\mathrm{T}}\!\!\!.$ (8)
The symbol $d_{n}(\phi,\psi,\mathbf{p})$ is referred to as the path difference of the signal propagation from the direction $(\phi,\psi)$ between the $n$-th element and the reference origin, which gives
@@
Without loss of generalization, the fRIS is typically deployed in the area where line-of-sight (LoS) paths dominate, so that the signal coverage can be extended. Consequently, the fRIS-related channels are modeled as LoS channels, which are determined by the steering vector of the fRIS. For the channel $\mathbf{G}$ from the BS to the fRIS, it can be modeled as
@@
in which $\sqrt{\zeta_{G}}$ denotes the path loss between the BS and the fRIS, while $\mathbf{a}_{r}(\phi_{r},\psi_{r},\mathbf{p})$ represents the receive steering vector of the fRIS with the signal coming from $\phi_{r}$ in azimuth direction and $\psi_{r}$ in elevate direction, which is in the form of (8). On the other hand, the ULA is equipped in the BS, so that the transmit steering vector $\mathbf{a}_{t}$ in the direction of $\phi_{t}$ can be written as
@@
where $\{\widetilde{d}_{m}=\frac{\lambda}{2}m\sin(\phi_{t}),~{}\forall m=0,\cdots,M-1\}$ and the inter-element spacing is half wavelength. Similarly, the channel $\mathbf{h}_{rc,k}^{\mathrm{H}}$ from the fRIS to the user $k$ is modeled as
@@
where $\mathbf{a}_{c,k}(\phi_{c,k},\psi_{c,k},\mathbf{p})$ denotes the transmit steering vector of the fRIS to user $k$ from azimuth direction $\phi_{c,k}$ and elevate direction $\psi_{c,k}$, which takes the form of (8). The path loss between the fRIS and the user $k$ is denoted as $\sqrt{\zeta_{k}}$. Accordingly, the overall reflective channel $\mathbf{H}_{rc}$ can be rewritten as
@@
where we denote
@@
$\boldsymbol{\Sigma}_{rc}$ $=\mathrm{Diag}\left(\begin{bmatrix}\sqrt{\zeta_{1}}&\cdots&\sqrt{\zeta_{K}}\end{bmatrix}\right).$ (14b)
@@ II-B Problem Formulation
In order to simultaneously achieve the sensing and the communication functions, a weighted sum of the sensing beampattern mismatch $\varepsilon_{r}$ of (4) and the communication symbol estimation MSE $\varepsilon_{c}$ of (7) is minimized by optimizing the transmit beamformer $\mathbf{W}$, the fRIS phase shift $\boldsymbol{\Theta}$, the fRIS element positions $\mathbf{p}$, the communication symbol estimator $\omega$ as well as the scaling factor $\beta$ of the sensing beampattern. Consequently, the problem can be formulated as
$\min_{\mathbf{W},\boldsymbol{\Theta},\mathbf{p},\omega,\beta}$ $~{}\varepsilon_{0}\!=\!\alpha\left(\sum^{I_{a}}_{i=1}\sum^{I_{e}}_{i^{\prime}=1}\left|\beta\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})-\mathcal{P}_{s}(\phi_{i},\psi_{i^{\prime}})\right|^{2}\right)$
@@
where $\alpha$ is the weighting coefficient indicating the bias of the sensing and the communication. The constraint (15b) represents that the transmit power should not exceed $P_{t}$, while in the constraint (15d), the modulus-one phase shift of the fRIS is required. The constraints (15e) and (15f) indicate that each RIS element ought to move within the pre-set planar region, and the inter-space among elements should be larger than $\Delta D$. It is challenging to solve the problem (15) owing to the two reasons. On the one hand, (15d) is a constant modulus constraint, while (15f) takes the non-convex form of “Norm $\geq$ Constant”. Besides, the objective function is highly complex, where the term of beampattern mismatch $\varepsilon_{r}$ in the objective function is high-order with respect to different variables.
@@ III Proposed Algorithm
In this section, an effective AM-based algorithm is developed to address the problem (15) by iteratively updating the variables, and we analyze the convergence and computational complexity of the proposed algorithm.
@@ III-A Reformulation of $\varepsilon_{r}$
Carefully inspecting $\varepsilon_{r}$ in (4), it can be observed that $\varepsilon_{r}$ is a quadratic function with respect to the covariance matrix $\mathbf{R}_{s}$, while $\mathbf{R}_{s}$ is also a quadratic function with respect to the variables, making $\varepsilon_{r}$ a quartic function with respect to the variables. To simplify the objective function in (15a), we reduce the order of $\varepsilon_{r}$ to be a quadratic function with respect to the variables.
Specifically, we first design an optimum covariance matrix $\widetilde{\mathbf{R}}_{s}$ that only takes the sensing into consideration. According to [35], the problem of designing $\widetilde{\mathbf{R}}_{s}$ can given as
@@
where the objective function requires the designed sensing beampattern to be similar to an ideal one, the constraint (16b) indicates that $\widetilde{\mathbf{R}}_{s}$ is semi-definite, the radiation power at the fRIS is $\|\mathbf{Gx}\|^{2}$ in the constraint (16c), and the last constraint requires $\widetilde{\mathbf{R}}_{s}$ to be rank-one since the ultimately designed $\mathbf{R}_{s}$ should satisfy $\mathbf{R}_{s}=\mathbf{v}\mathbf{v}^{\mathrm{H}}$. The problem (16) can be directly solved with SDR approach. To guarantee the rank-one constraint of $\widetilde{\mathbf{R}}_{s}$, various methods can be employed, such as Gaussian randomization [36] and penalty-based method [37].
With the obtained $\widetilde{\mathbf{R}}_{s}$ designed in (16), the beams can be effectively formed towards the targets. Considering that the designed $\widetilde{\mathbf{R}}_{s}$ is rank-one, the corresponding reflective signal at the fRIS associated with $\widetilde{\mathbf{R}}_{s}$ can be retrieved by the performing eigenvalue decomposition to $\widetilde{\mathbf{R}}_{s}$. This yields
@@
where $\mathbf{s}_{r}=\sqrt{\lambda_{s}}\mathbf{u}_{s}$, while $\lambda_{s}$ and $\mathbf{u}_{s}$ represent the maximum eigenvalue and the corresponding eigenvector of $\widetilde{\mathbf{R}}_{s}$, respectively. The factorized signal $\mathbf{s}_{r}$ represents an optimum and achievable reflective signal whose correlation matrix is $\widetilde{\mathbf{R}}_{s}$ producing directive beams towards the targets. In other words, $\mathbf{s}_{r}$ can be viewed as a desired reflective signal for sensing the targets.
Consequently, the sensing metric is reformulated as MSE between the practical reflective signal $\mathbf{v}$ at the fRIS and the desired reflective signal $\mathbf{s}_{r}$, i.e.
@@
It can be observed that the reformulated sensing metric $\widetilde{\varepsilon}_{r}$ becomes quadratic with respect to all the variables, which simplifies the subsequent optimization.
Ultimately, the problem (15) can be reformulated as
$\min_{\mathbf{W},\omega,\mathbf{p},\boldsymbol{\Theta}}$ $~{}\varepsilon=\alpha\left(\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\right)$
@@
However, the problem is still challenging owing to the non-convex constraints and objective function. In the subsequence, the problem (19) is solved by iteratively updating each variable with others being fixed. It is worth mentioning that $\mathbf{s}_{r}$ may change after the update of $\mathbf{p}$, but according to the Lyapunov stability [38], $\mathbf{s}_{r}$ is asymptotically stable over iterations due to the bounded $\mathbf{p}$ and the finite resources.
@@ III-B Optimization of Communication Symbol Estimator $\omega$
When $\mathbf{x}$, $\mathbf{p}$ and $\boldsymbol{\Theta}$ are fixed, the objective function in the problem (19) is a quadratic function with respect to $\omega$. After ignoring the terms that are unrelated to $\omega$, the problem can be simplified as
@@
The optimal $\omega^{\star}$ can therefore be derived by letting the derivative of $\varepsilon_{c}$ be zero, that is, $\frac{\partial\varepsilon_{c}}{\partial\omega}=0$. This yields a closed-form optimum:
@@ III-C Optimization of fRIS Phase Shift $\boldsymbol{\Theta}$
Given $\mathbf{W}$, $\mathbf{p}$ and $\omega$, the problem (19) can be reduced to a constant-modulus constrained quadratic minimization problem, which is given as
@@
The challenge of this problem mainly stems from the non-convex constant-modulus constraint. To address this task, one can first let $\boldsymbol{\theta}=\mathrm{diag}(\boldsymbol{\Theta})$ and employ the property of $\mathrm{tr}(\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{C}_{1}\boldsymbol{\Theta}\mathbf{C}_{2})=\boldsymbol{\theta}^{\mathrm{H}}(\mathbf{C}_{1}\odot\mathbf{C}_{2}^{\mathrm{T}})\boldsymbol{\theta}$ [21] to transform the problem (22) to a compact form:
@@
where
${\mathbf{A}}_{1}$ $\!=\!\alpha\left[\left(\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\right)\odot\mathbf{I}\right]$
$~{}~{}+\omega^{2}\left(1\!-\!\alpha\right)\!\left[\left(\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\right)\!\odot\!\left(\mathbf{H}_{rc}^{*}\mathbf{H}_{rc}^{\mathrm{T}}\right)\right],$ (24a)
${\mathbf{b}}_{1}$ $=\alpha\mathrm{diag}(\mathbf{G}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}})+\omega(1-\alpha)\mathrm{diag}(\mathbf{G}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{rc}^{\mathrm{H}}).$ (24b)
The nonconvex constant-modulus constraint (23b) remains challenging, where various approaches have been developed in the wide literature to effectively handle such a constraint, such as SDR [39] and manifold optimization [40]. Therefore, the details will not be elaborated upon here.
III-D Optimization of Transmit Beamformer $\mathbf{W}$
In this subsection, the optimization of the transmit beamformer $\mathbf{W}$ will be discussed. Viewing other variables as constant, the problem (19) can be reduced to
@@ III-D Optimization of Transmit Beamformer $\mathbf{W}$
$\min_{\mathbf{W}}$ $~{}\alpha\!\left(\!\|\mathbf{s}_{r}\!-\!\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{W}\mathbf{s}_{c}\|^{2}\right)\!\!+\!(1\!-\!\alpha)\!\left(\|\mathbf{s}_{c}\!-\!\omega\mathbf{H}_{c}\mathbf{W}\mathbf{s}_{c}\|^{2}\right)$ (25a)
$\mathrm{s.t.}$ $~{}\|\mathbf{W}\|^{2}\leq P_{t},$ (25b)
@@
where we denote $\mathbf{H}_{c}=\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}$ as the overall channel from the transmitter to the users. For the objective function (25a), the norm operations are expanded and thus it becomes
@@
$\alpha\!\left(\mathbf{s}_{c}^{\mathrm{H}}\mathbf{W}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{W}\mathbf{s}_{c}-2\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{W}\mathbf{s}_{c}\}\right)$
$+$ $(1\!-\!\alpha)\!\left(\omega^{2}\mathbf{s}_{c}^{\mathrm{H}}\mathbf{W}^{\mathrm{H}}\mathbf{H}_{c}^{\mathrm{H}}\mathbf{H}_{c}\mathbf{W}\mathbf{s}_{c}-2\omega\mathfrak{R}\{\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{c}\mathbf{W}\mathbf{s}_{c}\}\right).$ (26)
By leveraging the cyclic property of trace, the relationship of $\mathrm{tr}(\mathbf{X}^{\mathrm{H}}\mathbf{C}_{1}\mathbf{X}\mathbf{C}_{2})=\mathrm{vec}^{\mathrm{H}}(\mathbf{X})(\mathbf{C}_{2}^{\mathrm{T}}\otimes\mathbf{C}_{1})\mathrm{vec}(\mathbf{X})$ and $\mathrm{tr}(\mathbf{C}^{\mathrm{H}}\mathbf{X})=\mathrm{vec}^{\mathrm{H}}(\mathbf{C})\mathrm{vec}(\mathbf{X})$, the objective function can further be simplified as
$\mathbf{w}^{\mathrm{H}}\mathbf{A}_{2}\mathbf{w}-2\mathfrak{R}\{\mathbf{b}_{2}^{\mathrm{H}}\mathbf{w}\},$ (27)
in which
$\mathbf{w}$ $=\mathrm{vec}(\mathbf{W}),$ (28a)
$\mathbf{A}_{2}$ $\!=\!\alpha\!\left(\mathbf{s}_{c}^{*}\mathbf{s}_{c}^{\mathrm{T}}\!\otimes\!\mathbf{G}^{\mathrm{H}}\mathbf{G}\right)\!+\!(1\!-\!\alpha)\omega^{2}\!\left(\mathbf{s}_{c}^{*}\mathbf{s}_{c}^{\mathrm{T}}\!\otimes\!\mathbf{H}_{c}^{\mathrm{H}}\mathbf{H}_{c}\right),$ (28b)
@@
$\mathbf{b}_{2}$ $=\alpha\left(\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\mathbf{s}_{r}\mathbf{s}_{c}^{\mathrm{H}}\right)+(1\!-\!\alpha)\omega\left(\mathbf{H}_{c}^{\mathrm{H}}\mathbf{s}_{c}\mathbf{s}_{c}^{\mathrm{H}}\right).$ (28c)
It is worth noting that in (28b), $\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}$ has been simplified as $\mathbf{G}^{\mathrm{H}}\mathbf{G}$ since $\boldsymbol{\Theta}$ is a modulus-one diagonal matrix and $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}=\mathbf{I}$ always holds. After this transformation, it can be seen that the objective (27) is a convex quadratic function. As for the constraint (25b), it can also be re-expressed as the inner product of $\mathbf{w}$, that is, $\mathbf{w}^{\mathrm{H}}\mathbf{w}\leq P_{t}$. In this context, we consider full power usage in the transmitter, namely $\mathbf{w}^{\mathrm{H}}\mathbf{w}=P_{t}$. As a result, the problem (25) can be written as
$\min_{\mathbf{w}}$ $~{}\mathbf{w}^{\mathrm{H}}\mathbf{A}_{2}\mathbf{w}-2\mathfrak{R}\{\mathbf{b}_{2}^{\mathrm{H}}\mathbf{w}\}$ (29a)
$\mathrm{s.t.}$ $~{}\mathbf{w}^{\mathrm{H}}\mathbf{w}=P_{t},$ (29b)
which can be addressed by ALM [41, 42]. Specifically, we first employ the complex-real conversion by separating the real part and the imaginary part. The problem can thus be converted to its real form as
$\min_{\overline{\mathbf{w}}}$ $~{}\mathring{f}(\overline{\mathbf{w}})=\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{A}}_{2}\overline{\mathbf{w}}-2\overline{\mathbf{b}}_{2}^{\mathrm{T}}\overline{\mathbf{w}}$ (30a)
$\mathrm{s.t.}$ $~{}\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}=P_{t},$ (30b)
where
$\overline{\mathbf{w}}$ =[$=\begin{bmatrix}\mathfrak{R}\{\mathbf{w}\}\\ \mathfrak{I}\{\mathbf{w}\}\end{bmatrix},\overline{\mathbf{b}}_{2}=\begin{bmatrix}\mathfrak{R}\{\mathbf{b}_{2}\}\\ \mathfrak{I}\{\mathbf{b}_{2}\}\end{bmatrix},$ (31a)
$\overline{\mathbf{A}}_{2}$ =[$=\begin{bmatrix}\mathfrak{R}\{\mathbf{A}_{2}\}&-\mathfrak{I}\{\mathbf{A}_{2}\}\\ \mathfrak{I}\{\mathbf{A}_{2}\}&\mathfrak{R}\{\mathbf{A}_{2}\}\end{bmatrix}.$ (31b)
Accordingly, the augmented Lagrangian function of problem (30) can be written as
$\mathcal{L}\triangleq\mathring{f}(\overline{\mathbf{w}})+\mu\left(\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}-P_{t}\right)+\frac{\gamma}{2}\left(\overline{\mathbf{w}}^{\mathrm{T}}\overline{\mathbf{w}}-P_{t}\right)^{2},$ (32)
in which $\mu$ is the Lagrange multiplier and $\gamma$ denotes the penalty factor. This can be tackled by the following iteration procedure:
$\overline{\mathbf{w}}^{(t)}$ $=\arg\min_{\overline{\mathbf{w}}}\mathcal{L}(\overline{\mathbf{w}},\mu^{(t-1)}),$ (33a)

### v3

@@ II-A3 Channel Model
The steering vector of the fRIS is dependent on the positions of the elements $\mathbf{p}$, the azimuth angle $\phi$, and the elevation angle $\psi$. Specifically, the steering vector of the fRIS is expressed as [23]:
$\!\!\mathbf{a}(\phi,\psi,\mathbf{p})\!=\!\begin{bmatrix}e^{\jmath\frac{2\pi}{\lambda}d_{1}(\phi,\psi,\mathbf{p}_{1})}&\!\!\!\cdots\!\!\!&e^{\jmath\frac{2\pi}{\lambda}d_{N}(\phi,\psi,\mathbf{p}_{N})}\end{bmatrix}^{\mathrm{T}}\!\!\!,$ (7)
where $d_{n}(\phi,\psi,\mathbf{p})$ is referred to as the path difference of the signal propagation from the direction $(\phi,\psi)$ between the $n$-th element and the reference origin, given by
@@
Without loss of generality, the fRIS is considered deployed in the area where line-of-sight (LoS) paths dominate, such that the signal coverage can be extended. Consequently, the fRIS-related channels are modeled as LoS channels. The channel from the BS to the fRIS is modeled as
@@
in which $\sqrt{\zeta_{G}}$ denotes the path loss between the BS and the fRIS, while $\mathbf{a}_{r}(\phi_{r},\psi_{r},\mathbf{p})$ represents the receive steering vector of the fRIS with the signal coming from $\phi_{r}$ in azimuth and $\psi_{r}$ in elevation. On the other hand, as the ULA is equipped in the BS, the transmit steering vector in the direction of $\phi_{t}$ is written as
@@
where $\{\widetilde{d}_{m}=\frac{\lambda}{2}m\sin(\phi_{t}),~\forall m=0,\cdots,M-1\}$ and the inter-element spacing is half wavelength. Similarly, the channel from the fRIS to the $k$-th user is expressed as
@@
where $\mathbf{a}_{c,k}(\phi_{c,k},\psi_{c,k},\mathbf{p})$ denotes the steering vector of the fRIS to the $k$-th user from $\phi_{c,k}$ and $\psi_{c,k}$. The path loss between the fRIS and the $k$-th user is denoted as $\sqrt{\zeta_{k}}$. Accordingly, the overall downlink reflective channel $\mathbf{H}_{rc}$ is rewritten as
@@
where $\mathbf{A}_{rc}$ and $\mathbf{\Sigma}_{rc}$ are defined as
@@
[…]
Note: In practice, RIS element positions may deviate slightly from their designed values due to hardware imperfections. Therefore, deviation $\mathbf{p}=\overline{\mathbf{p}}+\Delta\mathbf{p}$ may introduce phase shifts offset in the steering vector $\mathbf{a}(\mathbf{p})=\boldsymbol{\Pi}(\Delta\mathbf{p})~\overline{\mathbf{a}}(\overline{\mathbf{p}})$. It can be analyzed that the effect on communication is negligible due to phase cancellation, while the sensing performance may experience distortion in the beampattern shape, which remains limited under small deviations. While the extension to a robust framework regarding position deviations is non-trivial, it is left to future work.
@@ II-B Problem Formulation
To simultaneously achieve sensing and communication functions, a weighted sum of the sensing beampattern mismatch $\varepsilon_{r}$ of (3) and the communication symbol estimation MSE $\varepsilon_{c}$ of (II-A2) is minimized by optimizing the transmit waveform $\mathbf{x}$, the fRIS phase shift $\boldsymbol{\Theta}$, the fRIS element positions $\mathbf{p}$, the communication symbol estimator $\omega$ and the scaling factor $\beta$ of the sensing beampattern. Consequently, the problem is formulated as
$\min_{\mathbf{x},\boldsymbol{\Theta},\mathbf{p},\omega,\beta}$ $~\varepsilon_{0}\!=\!\alpha\left(\sum^{I_{a}}_{i=1}\sum^{I_{e}}_{i^{\prime}=1}\left|\beta\mathcal{P}_{d}(\phi_{i},\psi_{i^{\prime}})-\mathcal{P}_{s}(\phi_{i},\psi_{i^{\prime}})\right|^{2}\right)$
@@
where $\alpha$ is the weighting coefficient, controlling the trade-off between sensing and communication. The constraint of (14b) indicates that the transmit power should not exceed $P_{t}$, while (14d) requires the modulus-one phase shifts of the fRIS. The constraints of (14e) and (14f) are introduced to model the physical limitations of the fRIS structure. Specifically, (14e) captures the physical boundary of the movement area, ensuring that all elements are confined in $\mathcal{A}$. Besides, (14f) enforces a minimum distance $\Delta D$ among elements, mitigating the electromagnetic mutual coupling effects and avoiding element collision.
The problem (14) is non-convex owing to the two reasons. On the one hand, (14d) and (14f) are non-convex. Besides, the objective function is complex, where the term of beampattern mismatch $\varepsilon_{r}$ in the objective function is high-order with respect to (w.r.t.) different variables, while the variables are coupled and not jointly convex.
@@ III Proposed Algorithm
In this section, an AM-based algorithm is developed to address (14). Given the non-convexity and NP-hardness of the problem, we solve it by iteratively updating the variables. Additionally, we analyze the convergence and the computational complexity of the proposed algorithm.
@@ III-A Reformulation of $\varepsilon_{r}$
Carefully inspecting $\varepsilon_{r}$ in (3), it is a quartic function of the optimization variables, as it is quadratic w.r.t. $\mathbf{R}_{s}$, which itself depends quadratically on the optimization variables. To simplify the objective function in (14a), we reduce the order of $\varepsilon_{r}$ to be a quadratic function w.r.t. the variables. First, we design a reference covariance matrix $\widetilde{\mathbf{R}}_{s}$ that only considers sensing. According to [16], the problem of designing $\widetilde{\mathbf{R}}_{s}$ can be given by
@@
where the objective function requires the designed sensing beampattern to be similar to an ideal one. In the constraint of (15c), the radiation power at the fRIS is $\|\mathbf{Gx}\|^{2}$, and the last constraint requires $\widetilde{\mathbf{R}}_{s}$ to be rank-one since $\mathbf{R}_{s}=\mathbf{v}\mathbf{v}^{\mathrm{H}}$. This problem can be directly solved via SDR with Gaussian randomization [21] and penalty-based method [52].
Remark 1: To handle (15d), SDR followed by Gaussian randomization can be adopted to construct a feasible rank-one solution. While a theoretical proof of tightness is generally intractable, this approach is widely adopted and has demonstrated strong empirical performance. For the penalty-based method, the rank-one constraint can be enforced via a penalty term $\big(\mathrm{tr}(\widetilde{\mathbf{R}}_{s})-\|\widetilde{\mathbf{R}}_{s}\|_{2}\big)$. As proven in [50], a rank-one solution can be obtained when the penalty factor is sufficiently large.
With the obtained $\widetilde{\mathbf{R}}_{s}$ in (15), the beams can be effectively formed towards the targets. Since $\widetilde{\mathbf{R}}_{s}$ is rank-one, a reference signal $\mathbf{s}_{r}$ at the fRIS associated with $\widetilde{\mathbf{R}}_{s}$ can be obtained by the eigenvalue decomposition of $\widetilde{\mathbf{R}}_{s}$, yielding
@@
where $\mathbf{s}_{r}=\sqrt{\lambda_{s}}\mathbf{u}_{s}$ with $\lambda_{s}$ and $\mathbf{u}_{s}$ being the maximum eigenvalue and the associated eigenvector, respectively. Here, $\mathbf{s}_{r}$ represents a desired and achievable reflective signal, which produces directive beams towards the targets.
Consequently, the sensing metric is reformulated as the mismatching between $\mathbf{v}$ and $\mathbf{s}_{r}$, i.e.,
@@
It is observed that $\widetilde{\varepsilon}_{r}$ becomes quadratic w.r.t. all variables, which simplifies the subsequent optimization.
Remark 2: The reformulation approximation may incur error from two sources: (i) the rank-one recovery of $\widetilde{\mathbf{R}}_{s}$ via decomposition, which is mainly from the potential suboptimality of Gaussian randomization, and (ii) the imperfect realization of $\mathbf{s}_{r}$ via $\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}$ under the joint ISAC design. The dominant error arises from the mismatch $\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}$, which is minimized. Although theoretical bounds are difficult to obtain, the approximation error remains trivial in practice.
Ultimately, (14) is reformulated as
$\min_{\mathbf{x},\omega,\mathbf{p},\boldsymbol{\Theta}}$ $~\varepsilon=\alpha\left(\|\mathbf{s}_{r}-\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\right)$
@@
Subsequently, we solve (18) by iteratively updating the variables.
@@ III-B Optimization of Communication Symbol Estimator $\omega$
Concerning $\omega$, (18a) is a quadratic function. After ignoring the terms that are unrelated to $\omega$, (18) is simplified as
@@
The optimal $\omega^{\star}$ can be derived by setting the derivative of $\varepsilon_{c}$ to be zero, that is, $\frac{\partial\varepsilon_{c}}{\partial\omega}=0$, yielding the closed-form solution:
@@ III-C Optimization of fRIS Phase Shift $\boldsymbol{\Theta}$
Given $\mathbf{x}$, $\mathbf{p}$, and $\omega$, (18) can be reduced to a modulus-one constrained quadratic minimization problem, given by
@@
To address (21), we first define $\boldsymbol{\theta}=\mathrm{diag}(\boldsymbol{\Theta})$ and employ the property of $\mathrm{tr}(\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{C}_{1}\boldsymbol{\Theta}\mathbf{C}_{2})=\boldsymbol{\theta}^{\mathrm{H}}(\mathbf{C}_{1}\odot\mathbf{C}_{2}^{\mathrm{T}})\boldsymbol{\theta}$ [48] to transform (21) to a compact form:
@@
where ${\mathbf{b}}_{1}\!=\!\alpha\mathrm{diag}(\mathbf{G}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}})\!+\omega(1\!-\!\alpha)\mathrm{diag}(\mathbf{G}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{rc}^{\mathrm{H}})$ and ${\mathbf{A}}_{1}\!=\!\alpha\left[\left(\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\right)\odot\mathbf{I}\right]+\omega^{2}\left(1\!-\!\alpha\right)\!\left[\left(\mathbf{G}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\right)\!\odot\!\left(\mathbf{H}_{rc}^{*}\mathbf{H}_{rc}^{\mathrm{T}}\right)\right]$. To address (22b), various approaches can be adopted, such as SDR [42] and manifold optimization [49]. While SDR produces suboptimal solutions due to rank-one relaxation, it ensures a worst-case $\pi/4$-approximation. Manifold optimization performs gradient descent over the complex circle manifold, transforming the problem into an unconstrained search. Despite inherent non-convexity, both methods achieve high-quality local optima and demonstrate strong empirical performance.
III-D Optimization of Transmit Waveform $\mathbf{x}$
In this subsection, the optimization of the transmit waveform $\mathbf{x}$ is discussed. First, (18) w.r.t. $\mathbf{x}$ is simplified as
@@ III-D Optimization of Transmit Waveform $\mathbf{x}$
$\min_{\mathbf{x}}$ $~\alpha\!\left(\!\|\mathbf{s}_{r}\!-\!\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\|^{2}\right)\!\!+\!(1\!-\!\alpha)\!\left(\|\mathbf{s}_{c}\!-\!\omega\mathbf{H}_{c}\mathbf{x}\|^{2}\right)$ (23a)
$\mathrm{s.t.}$ $~\|\mathbf{x}\|^{2}\leq P_{t},$ (23b)
@@
where $\mathbf{H}_{c}=\mathbf{H}_{rc}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\in\mathbb{C}^{K\times M}$ is denoted as the overall channel from the transmitter to the users. For (23a), the squared norm is expanded, and the constant terms are omitted, leading to the equivalent expression as
@@
$\alpha\!\left(\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}-2\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\}\right)$
$\!\!\!+(1\!-\!\alpha)\!\left(\omega^{2}\mathbf{x}^{\mathrm{H}}\mathbf{H}_{c}^{\mathrm{H}}\mathbf{H}_{c}\mathbf{x}-2\omega\mathfrak{R}\{\mathbf{s}_{c}^{\mathrm{H}}\mathbf{H}_{c}\mathbf{x}\}\right).$ (24)
The objective function of (III-D) is compactly rewritten as
$\mathbf{x}^{\mathrm{H}}\mathbf{A}_{2}\mathbf{x}-2\mathfrak{R}\{\mathbf{b}_{2}^{\mathrm{H}}\mathbf{x}\}$ (25)
by denoting
@@
$\mathbf{A}_{2}$ $\!=\!\alpha\!\left(\!\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\!\right)\!+\!(1\!-\!\alpha)\omega^{2}\!\left(\!\mathbf{H}_{c}^{\mathrm{H}}\mathbf{H}_{c}\right)\!,$ (26a)